The Taipei 101 building in Taipei, Taiwan, is
1670 ft tall. How long would it take an object to fall from
the top?

Answer by stanbon(75887)

The Taipei 101 building in Taipei, Taiwan, is
1670 ft tall. How long would it take an object to fall from
the top?
-------
Solve -16t^2 = -1670
t^2 = 1670/16
t = 10.22 seconds
----------------
Cheers,
Stan H.
--------------